OpenStudy (nicoledonley):

If x^2-y^2=12 and x-y=4, what is the value o x^2+2xy+y^2?

OpenStudy (mww):

hint: factorise x^2 - y^2 and x^2 + 2xy + y^2 separately

OpenStudy (nicoledonley):

So (x+y)(x-y)=12

OpenStudy (jiteshmeghwal9):

(x+y)(x-y)=12.....(1) (x-y)=4....(2) use equation 1 & 2 to find (x+y), then \(x^2+2xy+y^2=(x+y)^2=?\)